Assume a DVC and an IAC currently have real per capita outputs of $1,000 and $10,000, respectively. Instructions:

If both nations have a 5 percent increase in their real per capita outputs, by how much will the per capita output gap change?
Answer #1

Output Gap between DVC and IAC = 10000-1000 = 9000

After 5% increase,Real per capita output of DVC = 1050 and Real per capita output of IAC = 10500

Output Gap = 10500-1050 = 9450

Output Gap change = 9450-9000 = 450